# GRS: Generating Robotic Simulation Tasks from Real-World Images

### v1

@@ IV-B Robot Task Generation
We designed our simulation generation experiment to assess GRS’s capacity to produce working simulators for robot policies. Simulations use the CLIPort [30] task framework and we base our generation pipeline on the prompts used in GenSim [9] with minor modifications to indicate VLMs should use an input image. We evaluate task generation systems by assessing how well an oracle policy can complete the generated simulation tasks. Specifically, we use three runs of executing GenSim’s oracle policy and average the results over these runs.
We compare our method (ours) to three ablations: 1) removing the router and only fixing the simulation (no router), 2) removing the tests and only generating the simulation once (no fix) and, 3) removing the image input and using an LLM only for generating the simulation once (LLM) The last ablation is the closest to the original GenSim [9]. The simulation generation only uses GPT4o given its superior performance in object correspondence. For each of our 10 scenes we run 30 generations and allow up to 10 attempts to fix the code (either simulation or tests) in each run. The same process is followed for all ablations, where no router is limited to 10 test fixes and ours is limited to a combined total of 10 fixes of either the simulation or tests.
TABLE II: Rewards obtained on generated tasks.
@@
Simulations generated by our method enable the oracle policy to obtain higher rewards than alternatives, see Table II. We report average rewards over the oracle policy executions for each generated simulation over all 10 scenes. As our interest is using these environments for robotic training we exclude cases with runtime errors. All results are normalized where a reward of 1 indicates completion of all goals in the task. We find our method yields simulations that are effective for policy execution, and that removing test feedback (no fix and LLM) substantially lowers policy execution success. Figure 3 presents qualitative results of generated simulations from our method.
@@
We also examined the simulation and test fixing behavior to understand the router’s behavior. The router performs 0.52 fixes to tests on average across tasks, indicating there is a frequent need to repair the initially generated tests. In addition, the router performs 1.08 fewer simulation fixes compared to no router (5.81 vs 6.89), indicating the router is more efficient when making changes, making net 0.56 fewer total changes per generation. This supports the claim that the router enables more efficient automation of simulation generation.
IV-C Qualitative Analysis of Code Generation
@@
To better understand GRS’s behavior, we qualitatively discuss the behavior of the router and the changes made when fixing the simulations and tests.
@@
Router. The router demonstrates a useful behavior of choosing to change the tests when the test feedback is too sparse to diagnose why the oracle fails. When errors occurred, the router appropriately parsed error feedback to recognize the need to fix tests when there were problems with missing imports or cases where the test misused simulation objects, e.g., assuming an object was a list when it was not.
@@
Simulation Fixing. Often the simulation task was too complex to be successfully executed by the oracle. We observed different behaviors to address this problem: 1) simplifying the simulation by reducing the number of objects used (thus simplifying the task goals), 2) increasing the maximum number of steps the oracle could take before terminating an attempt, or 3) increasing the size of target areas for placing objects. These fixes were typically in response to running tests where the oracle achieved only partial success on a task. This demonstrates that our proposed system is able to correctly parse nuanced results where no explicit error is raised beyond reporting that the oracle obtained a low reward. However, on occasion this would cause divergences from our intended outcome, where the fix would remove the provided object spatial locations and only retain the assets.
@@
Test Fixing. During test execution the oracle could fail the task but not produce meaningful errors due to the absence of simulator runtime errors. In response, the test fixes would add diagnostics about the oracle execution. These added monitoring for crucial performance indicators, including step count, intermediate goal achievement, and reward accumulation, thereby providing feedback signals throughout the testing process. Moreover the test fixes extend beyond mere performance tracking, addressing fundamental issues such as erroneous interaction with simulation environment components and goal misinterpretation. The test fixes would also focus on validating the correct initialization and reset functionality of the simulation environment. These measures ensure proper goal setup and the generation of valid observational data. Note that these improvements were achieved while adhering to our principle of using generic prompts, which were designed solely to guide the system in testing oracle success. This underscores the effectiveness of our refined test protocols in enhancing system evaluation without altering the core testing paradigm. Further, this approach has promise for improving with better baseline language models, requiring no modifications to our framework.
@@
Common Failure Cases. During the experiments we discovered simulation code would sometimes (roughly 30 times in 1200 generations) include while loops that could potentially fail to terminate, e.g., looping indefinitely when trying to find a valid position to place assets in a constrained area. This behaviour would not always lead to failure, as most of the time the behaviour would be correct. Further, the non-terminating loop condition meant that it could cause executing the simulation to never exit, hanging the generation process. This problem could be addressed by explicitly prompting the code generation and fixing processes to avoid using these kinds of loops, encouraging avoiding non-termination problems, or implementing timeout behavior in the baseline generation and testing framework.
@@
During the fixing process the LLM would occasionally misdiagnose the cause of test failures, leading to implementing new code as workarounds. In simple cases the LLM would write geometric manipulation and other primitive functions instead of using the provided API, e.g., converting 3D rotations between quaternions and Euler angle representations, sampling from a probability distribution, inverting poses, etc. In rare cases this led to re-implementing the underlying reward functions or oracle agent in their entirety. Another failure includes the LLM mocking out the simulation environment creation or key aspects of the simulation behavior, bypassing the desired behaviour all together. We anticipate improvements in LLM capabilities will yield better error diagnosis and code fixing over time.
IV-D Scene-level Extension
@@
To demonstrate the flexibility of our approach, we extended our pipeline to generate scene-level tasks using a larger asset repository of approximately 150,000 assets from Objaverse [31]. Starting from a single RGB observation, we reconstruct the background of the scene by employing background estimation, fitting an MLP to estimate the SDF of the background surface, and then applying the marching cubes algorithm to generate the background mesh, following Dogaru et al. [32]. Once the background is reconstructed, we use the GRS real-to-sim workflow—combining object segmentation with VLM-based object matching—to construct the 3D task environment, see Figure 4. This work is a preliminary step towards generating more complex scene-level tasks, where we leave further exploration of this direction for future work.
@@
V Conclusions
In this work, we presented GRS, a novel system for Generating Robotic Simulation tasks from real-world observations. Our approach seamlessly integrates scene understanding, asset population, and task generation to address the challenge of real-to-sim. We introduced a robust scene acquisition system leveraging SAM2 [15] and vision-language models, integrated an LLM-based task generation framework, and proposed an innovative iteration technique using a dual-generation process and router system. Our experimental results demonstrate GRS’s effectiveness in achieving accurate real-to-sim translation from a single RGB-D observation, showcasing the power of VLMs in bridging the gap between real-world scenes and simulated tasks. We also presented another possible application of our approach with much wider asset pools and diversity.
The implications of this work extend beyond robotics simulation, potentially accelerating research and development in robotic manipulation, virtual and augmented reality, autonomous systems training, and computer vision. While our current implementation shows promising results, future work could focus on handling more complex scenes, improving asset matching scalability, integrating physics-based reasoning, and exploring transfer learning techniques. As we continue to refine and expand this approach, we anticipate that GRS will play a crucial role in bridging the gap between real-world observations and simulated environments, ultimately leading to more capable and versatile robotic systems.
ACKNOWLEDGMENT
We would like to thank Balakumar Sundaralingam for helping us with our robotics system.

### v2

@@ 4.2 Robot Task Generation
We designed our simulation generation experiment to assess GRS’s capacity to produce working simulators for robot policies. Simulations use the CLIPort [31] task framework and we base our generation pipeline on the prompts used in GenSim [36] with minor modifications to indicate VLMs should use an input image. GenSim prompts provide few-shot examples for task definitions and simulation code, along with API definitions that include a template base class for simulation code. We evaluated task generation systems by assessing how well an oracle policy can complete the generated simulation tasks. For evaluation scoring we used three runs of executing GenSim’s oracle policy and averaged the results over these runs.
We compared our method (ours) to three ablations: 1) removing the router and only fixing the simulation (no router), 2) removing the tests and only generating the simulation once (no fix) and, 3) removing the image input and using an LLM only for generating the simulation once (LLM) The last ablation is the closest to the original GenSim [36]. For simulation generation we only evaluated GPT4o given its superior performance in object correspondence. For each of our 10 scenes we ran 30 generations and allowed up to 10 attempts to fix the code (either simulation or tests) in each run. The same process was followed for all ablations, where no router was limited to 10 test fixes and ours was limited to a combined total of 10 fixes of either the simulation or tests.
Table 2: Rewards obtained on generated tasks.
@@
Simulations generated by our method enabled the oracle policy to obtain higher rewards than alternatives, see Table 2. We report average rewards over the oracle policy executions for each generated simulation over all 10 scenes. As our interest is using these environments for robotic training we exclude cases with runtime errors. All results are normalized where a reward of 1 indicates completion of all goals in the task. We found our method yielded simulations that were effective for policy execution, and that removing test feedback (no fix and LLM) substantially lowered policy execution success. Figure 3 presents qualitative results of simulations generated by our method.
@@
We also examined the simulation and test fixing behavior to understand the router’s behavior. The router performed 0.52 fixes to tests on average across tasks, indicating there was a frequent need to repair the initially generated tests. In addition, the router performed 1.08 fewer simulation fixes compared to no router (5.81 vs 6.89), indicating the router was more efficient when making changes, making net 0.56 fewer total changes per generation. This supports the claim that the router enabled more efficient automation of simulation generation.
4.3 Qualitative Analysis of Code Generation
@@
To better understand GRS’s behavior, we qualitatively examined the behavior of the router and the changes made when fixing the simulations and tests.
@@
Router. The router demonstrated a useful behavior of choosing to change the tests when the test feedback is too sparse to diagnose why the oracle fails. When errors occurred, the router appropriately parsed error feedback to recognize the need to fix tests when there were problems with missing imports or cases where the test misused simulation objects, e.g., assuming an object was a list when it was not.
@@
Simulation Fixing. Often the simulation task was too complex to be successfully executed by the oracle. We observed different behaviors to address this problem: 1) simplifying the simulation by reducing the number of objects used and thus simplifying the task goals, 2) increasing the maximum number of steps the oracle could take before terminating an attempt, or 3) increasing the size of target areas for placing objects. These fixes were typically in response to running tests where the oracle achieved only partial success on a task. This demonstrates that our proposed system was able to correctly parse nuanced results where no explicit error is raised beyond reporting that the oracle obtained a low reward. However, on occasion this diverge from our intended outcome, where the fix removed the provided object spatial locations and only retained the assets.
@@
Test Fixing. During test execution the oracle could fail the task but not produce meaningful errors due to the absence of simulator runtime errors. In response, the test fixes added diagnostics about the oracle execution. These monitored key performance indicators, including step count, intermediate goal achievement, and reward accumulation, thereby providing feedback signals throughout the testing process. Moreover the test fixes extended beyond mere performance tracking, addressing fundamental issues such as erroneous interaction with simulation environment components and goal misinterpretation. The test fixes also focused on validating the correct initialization and reset functionality of the simulation environment. These measures ensured proper goal setup and the generation of valid observational data. Note that these improvements were achieved while adhering to our principle of using generic prompts, which were designed solely to guide the system in testing oracle success. This underscors the effectiveness of our refined test protocols in enhancing system evaluation without altering the core testing paradigm. Further, this approach has promise for improving with better baseline language models, requiring no modifications to our framework.
@@
Common Failure Cases. During the experiments we discovered simulation code had rarely (roughly 30 times in 1200 generations) included while loops that could fail to terminate, e.g., looping indefinitely when trying to find a valid position to place assets in a constrained area. This behavior did not always cause to failure or appear during testing, as most of the time the behavior would be correct. The non-terminating loop condition meant the simulation execution would never exit, hanging the generation process. This problem could be addressed by explicitly prompting to avoid using these kinds of loops and non-terminating behavior, or implementing timeout behavior in the baseline generation and testing framework.
@@
During the fixing process the LLM sometimes misdiagnosed the cause of test failures and implemented new code as workarounds. In simple cases the LLM would write geometric manipulation and other primitive functions instead of using the provided API, e.g., converting 3D rotations between quaternions and Euler angle representations, sampling from a probability distribution, inverting poses, etc. In rare cases this led to re-implementing the underlying reward functions or oracle agent in their entirety. Sometimes the LLM would mock out the simulation environment creation or key aspects of the simulation behavior, bypassing the desired behavior altogether. We anticipate improvements in LLM capabilities will yield better error diagnosis and code fixing over time, potentially addressing these limitations without requiring changes to our framework.
4.4 Scene-level Extension
While our primary experiments focused on table-top scenes with a constrained asset set, we also tested our approach’s scalability to more complex environments. We extended our pipeline to process scenes using approximately 150,000 assets from Objaverse [11], significantly increasing the asset database size.
@@ 4.4 Scene-level Extension
Starting from a single RGB observation, we reconstructed the background of the scene by employing background estimation, fitting an MLP to estimate the SDF of the background surface, and then applying the marching cubes algorithm to generate the background mesh, following Dogaru et al. [12]. Once the background was reconstructed, we used GRS’s real-to-sim workflow—combining object segmentation with VLM-based object matching—to construct the 3D task environment, see Figure 4. The same object correspondence process described in Section 3.1 was applied, but with the much larger Objaverse repository as the source of potential matching assets.
The results, as shown in Figure 4, demonstrate that our approach can successfully scale to more complex scenes and significantly larger asset databases. This extension represents a preliminary step towards generating more complex scene-level tasks, where we leave further development of this direction to future work, as discussed in the following section.
@@
4.5 Limitations and Future Work
Building on our experimental results and the scene-level extension, we identify several limitations and opportunities for future work. Our experiments followed GenSim in assuming basic physics properties for scene assets. Future asset databases could provide richer metadata including mass, inertia, friction properties, and stiffness localized to asset parts. Our method readily extends to retrieving such assets for appropriately equipped simulators. More ambitiously, foundation models could be used to reason about physics to design tasks targeted to specific objects’ physical properties. For the scene-level extension, further work is needed to generate semantically meaningful tasks for complex environments by extending our router to handle increased complexity and spatial relationships in full scenes beyond tabletop arrangements.
Future work could also explore sim-to-real training and transfer learning techniques. Integrating GRS with sim-to-real processes could leverage the router to refine automatically revise the simulation and tests to close the sim-to-real gap. Using differentiable physics engines like Newton [20] could accelerate agent training through the simulator for reinforcement learning. Combined with our router, this could streamline the sim-to-real pipeline by automating both simulation generation and policy training.
5 Conclusions
In this work, we presented GRS, a system for generating robotic simulation tasks from real-world observations. Our approach integrates scene understanding with VLMs, asset matching, and task generation through an innovative dual-generation process and router system. Experiments demonstrate effective real-to-sim translation from single RGB-D observations and scalability to more complex scenes.
The implications of this work extend beyond robotics to game development, where GRS can automate realistic environment creation, generate solvable game tasks, and reduce development cycles for applications requiring realistic physics interactions. This approach is especially valuable for serious games and educational applications where real-world fidelity affects learning outcomes. Future work could address more complex scenes, improve asset matching, integrate physics-based reasoning, and explore transfer learning techniques as we continue developing digital twins across application domains.